A truck pulls two trailers, coupled together by a strong trailer hitch. It starts from rest and accelerates to 0.31 m / s2 forward on a straight road. The mass of the truck is 5325 kg, the mass of the first trailer is 28,205 kg, and the mass of the second trailer is 22,045 kg. Calculates the strength of the tension on the hitch between the truck and the first trailer, and the tension between the first and second trailer.
1
Expert's answer
2021-05-05T17:55:01-0400

To be given in question

Acceleration (a)=0.31"m\/sec^2"

Mass of truck "(m_{T})=5325kg"

Mass of first trailer "({m_{tr1}})=28205kg"

Mass of second trailer"(m_{tr2})=22045kg"

To be asked in question

Tansion T1=?

Tansion T2=?

We know that according question



System mass (m)="m_T+m_{tr1}+m_{tr2}"

Put value

m=5325+28205+22045=55575kg

T=ma

Put value

System net tension


"T =55575\times0.31=17228.25N"

"T_{2}=m_{tr2}\times a"

"T_{2}=22045\times0.31=6834.95N"



"T_{1}=(m_{tr1}+m_{tr2})a"

"T_{1}=(28205+22045)\times0.31=15577.5N"